How many grams of H2O are formed if 44.4 g of C5H12 is mixed with 104 g of O2?
Answer #1

Ansd :

The balanced molecular equation is given as :

C5H12 + 8O2 = 5CO2 + 6H2O

Number of mol og C5H12 = mass / molar mass

= 44.4 g / 72.14878 g/mol = 0.615 mol

Mol O2 = 104 g / 31.999 g/mol = 3.25 mol

each mol C5H12 needs 8 mol O2

So 0.615 mol C5H12 will need : 0.615 mol x 8 = 4.92 mol O2

we have only 3.25 mol oxygen so it is the limiting reagent

8 mol O2 forms 6 mol H2O

So 3.25 mol O2 will form : ( 3.25 mol x 6 mol) / 8 mol = 2.44 mol H2O

Mass of H2O formed = mol x molar mass

= 2.44 mol x 18.01528 g/mol

= 43.9 grams